OpenStudy (anonymous):

l Medals 0 For the right triangle shown, the lengths of two sides are given. Find the third side. Leave your answers in simplified, radical form

OpenStudy (anonymous):



OpenStudy (anonymous):

use Pythagorus theorem. c^2=a^2+b^2

OpenStudy (anonymous):

a = 40, b = a0, c = 41

OpenStudy (anonymous):

there was some extra info

OpenStudy (anonymous):

b=a0??

OpenStudy (anonymous):

or is b the one you want to find?

OpenStudy (anonymous):

the ao means blank........ and yes i do want to find b

OpenStudy (anonymous):

41^2=40^2+b^2 41^2-40^2=b^2 sqrt(41^2-40^2)=b b=sqrt(41^2-40^2) Can you solve this from here?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[b=\sqrt(41^2-40^2)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know what to do next?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

you have to get the square of 41 and 40 first before subtracting.

OpenStudy (anonymous):

ok can u help

OpenStudy (anonymous):

sqrt(1681-1600) sqrt(81) 9

OpenStudy (anonymous):

all right got it. can you answer this one Create the equation to be used to find the missing lengths. Do not solve the equation